OpenStudy (anonymous):

f(x)=12/x(x+4) Find the domain and range,

OpenStudy (anonymous):

domain R numver set -{0,-4} range (0, infinity)

OpenStudy (anonymous):

@raheen the answer in the book is that y is all real numbers except -1 and -2... we have no idea how to do that.

OpenStudy (anonymous):

ok , I think I read it in different way than what you mean, let me check that

OpenStudy (anonymous):

Its calculus, am i in the wrong room?

OpenStudy (anonymous):

no you are in the right room

OpenStudy (anonymous):

i'll assume f(x)=12/(x(x+4)) domain=all values except values that make x(x+4)=0 so domain is all real numbers x except for x=0 and x=-4 range is all real numbers

OpenStudy (anonymous):

f(x) = y when you plug in -1 then -1 = x So you can test from the points 0, 1, 2, 3, etc...

OpenStudy (anonymous):

range always positive real numbers==> (0,+infinity)

OpenStudy (anonymous):

edarosa can you write the problem using the drawing board

OpenStudy (anonymous):